OpenStudy (anonymous):

Manuel will rent a car for the weekend. He can choose one of two plans. The first plan has an initial fee of $ 75 and costs an additional $ 0.10 per mile driven. The second plan has no initial fee but costs $ 0.60 per mile driven. How many miles would Manuel need to drive for the two plans to cost the same

OpenStudy (anonymous):

@k_lynn

OpenStudy (welshfella):

let x be the number of miles you need to set up an equation in x and solve it

OpenStudy (welshfella):

First plan Cost = 75 + 0.x 2nd plan Cost = 0.6x for the same cost: 75 + 0.1x = 0.6 x find x